Selected financial information about Backpacking Resources, Inc., is as follows:

Total assets $10,000,000
Total liabilities $5,000,000
Total preferred stock $1,000,000
Total annual preferred stock dividends $80,000
Net profits after tax $500,000
Number of shares of common stock outstanding 250,000 shares
Current market price of common stock $58.00 a share
Annual common stock dividends $4.06 a share

Using the company's financial information, compute the following:

  1. Dividend yield. Round the answer to two decimal places.
      %
  2. Book value per share. Round your answer to the nearest cent.
    $    per share
  3. Earnings per share. Round your answer to the nearest cent.
    $    per share
  4. P/E ratio. Round the answer to two decimal places.
      times
Answer #1

a) Dividend yield =4.06/58 =7.00%

b) Book value per share:

=(Total assets-Total liabilities-Preferred capital)/Outstanding shares

=(10000000-5000000-1000000)/250000

=16.00

c) EPS:

=(Net profit after tax-Preferred dividend)/No of equity shares

=(500000-80000)/250000

=1.68

d) P/E =Current price/EPS =58/1.68 =34.52